OpenStudy (anonymous):

indicate the equation of the given line in standard form. The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (mathmate):

The given line is in the point-slope form, passing through the point (3,4). If we replace (3,4) by (1,-2), that will be the required equation of the new line after converting it to standard form.

OpenStudy (anonymous):

so replace (3,4) with (1,-2)? then solve?

OpenStudy (mathmate):

Right!

OpenStudy (mathmate):

A line in point slope form has the equation: \[(y-y_1=m(x-x_1)\] where m is the slope, and \[(x_1,y_1)\] is the point though which the line passes.

OpenStudy (anonymous):

so y+2=2/3(x-1)?

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Yes, the first step is correct. Well done! Now solve for y to put it in standard form.

OpenStudy (anonymous):

how do i do that?!!!!

OpenStudy (mathmate):

Start with \[y+2=\frac{2}{3}(x-1)\] can you expand the right-hand side?